Define a bijective function. Explain with reasons whether the following functions are bijective or not. Find also the inverse of each of the functions.

             i.    f(x) = 4x+2, A=set of real numbers

             ii.   f(x) = 3+ 1/x, A=set of non zero real numbers

             iii.  f(x) = (2x+3) mod7, A=N7


1
Expert's answer
2021-09-21T11:39:46-0400

Solution:

A bijective function has no unpaired elements and satisfies both injective (one-to-one) and surjective (onto) mapping of a set A to a set B. Thus, bijective functions satisfy injective as well as surjective function properties and have both conditions to be true. In mathematical terms, let f: A → B is a function; then, f will be bijective if every element ‘b’ in the co-domain B, has exactly one element ‘a’ in the domain A, such that f (a) =b.

i.   f(x) = 4x+2, A=set of real numbers

One-to-one:

"f(x_1)=f(x_2)
\\\Rightarrow 4x_1+2=4x_2+2
\\\Rightarrow 4x_1=4x_2
\\\Rightarrow x_1=x_2"

Thus, it is one-to-one.

Onto:

"f(y)=x
\\\Rightarrow 4y+2=x
\\\Rightarrow y=\dfrac{x-2}4"

Now, for all y in R, there exists "\dfrac{x-2}4" in R such that

"f(\dfrac{x-2}4)=x"

Thus, it is onto.

Hence, it is bijective.

"f^{-1}(y)=\dfrac{y-2}4 \forall y\in A"

          ii.  f(x) = 3+ 1/x, A=set of non zero real numbers

One-to-one:

"f(x_1)=f(x_2)
\\\Rightarrow 3+\dfrac1{x_1}=3+\dfrac1{x_2}
\\\Rightarrow \dfrac1{x_1}=\dfrac1{x_2}
\\\Rightarrow x_1=x_2"

Thus, it is one-to-one.

Onto:

f is not onto, because "3 \in A" (codomain) has no preimage in A (domain).

Hence, it is not onto. Thus, not bijective.

So, we cannot find "f^{-1}"


             iii. Given

"f: \mathrm{A} \rightarrow A 
\\ defined\ by\ f(x)=(2 x+3)(\bmod 7)
\\
where \quad \mathrm{A}=\mathrm{N}_{7}=\{0,1,2,3,4,5,6\}"

We have "\quad f(0)=3, f(1)=5, f(2)=0, f(3)=2, f(4)=4, f(5)=6, f(6)=1"

Then f is one-one and onto.

"\Rightarrow" f is bijection.

"\therefore f^{-1}=\{0,1,2,3,4,5,6\}"